3. The body weighing 2.5kg is hung on the hook of a dynamometer. Find: a) The weight of the body when the body and the dynamometer are at rest. b) Body weight when together with the dynamometer they climb up with an acceleration of 5m / s²


1
Expert's answer
2021-02-08T18:42:43-0500

a)

"W=mg=2.5\ kg\cdot9.8\ \dfrac{m}{s^2}=24.5\ N."

b)

"N-mg=ma,""N=m(g+a)=2.5\ kg\cdot(9.8\ \dfrac{m}{s^2}+5.0\ \dfrac{m}{s^2})=37\ N."